OBSERVATION OF VINOXY (CH_{2}CHO) RADICAL IN OZONOLYSIS REACTIONS OF 2-BUTENES USING CAVITY RING-DOWN SPECTROSCOPY

\underline{JINGSONG ZHANG} and LIMING WANG, Department of Chemistry and Air Pollution Research Center, University of California, Riverside, CA 92521.

Ozonolysis reactions of alkenes are important oxidation pathways of alkenes in the atmosphere, and they are also significant sources of tropospheric hydroxyl radicals. In this work, ozone reactions with trans- and cis-2-butene are studied using cavity ring-down spectroscopy (CRDS). Vinoxy (CH_{2}CHO) radical, a proposed co-product of OH from dissociation of Criegee intermediates following the primary ozonolysis of 2-butenes, is observed. The vinoxy formation is found to decrease drastically with increasing pressure. These results, along with quantum chemistry studies, suggest that the Criegee intermediate syn-CH_{3}CHOO, produced from dissociation of the 2-butene primary ozonides, could isomerize to CH_{2}CHOOH and then dissociate into the CH_{2}CHO + OH products.